# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

### v2

@@ Abstract
Steering cooperative multi-agent reinforcement learning (MARL) towards desired outcomes is challenging, particularly when the global guidance from a human on the whole multi-agent system is impractical in a large-scale MARL. On the other hand, designing external mechanisms (e.g., intrinsic rewards and human feedback) to coordinate agents mostly relies on empirical studies, lacking a easy-to-use research tool. In this work, we employ multi-agent influence diagrams (MAIDs) as a graphical framework to address the above issues. First, we introduce the concept of MARL interaction paradigms, using MAIDs to analyze and visualize both unguided self-organization and global guidance mechanisms in MARL. Then, we design a new MARL interaction paradigm, referred to as the targeted intervention paradigm that is applied to only a single targeted agent, so the problem of global guidance can be mitigated. In our implementation, we introduce a causal inference technique—referred to as Pre-Strategy Intervention (PSI)—to realize the targeted intervention paradigm. Since MAIDs can be regarded as a special class of causal diagrams, a composite desired outcome that integrates the primary task goal and an additional desired outcome can be achieved by maximizing the corresponding causal effect through the PSI. Moreover, the bundled relevance graph analysis of MAIDs provides a tool to identify whether an MARL learning paradigm is workable under the design of an MARL interaction paradigm. In experiments, we demonstrate the effectiveness of our proposed targeted intervention, and verify the result of relevance graph analysis.
@@ 1 Introduction
These MARL interaction paradigms cannot be fully captured by the conventional MARL learning paradigms (i.e., independent learning and centralized training).
@@ 2 Multi-Agent Influence Diagrams (MAIDs)
One way to figure out the above issue is making these two agents’ decision rules matched (koller2003multi, ), e.g. choosing both agents’ decision rules together, analogous to centralized training in multi-agent reinforcement learning (oliehoek2008optimal, ).
@@ 3 Multi-Agent Reinforcement Learning Interaction Paradigms
We introduce the term MARL interaction paradigms to describe the ways in which agents’ interactions are structured—ranging from unguided self-organization to externally guided mechanisms such as intrinsic rewards or human feedback. Specifically, we first formalize two existing MARL interaction paradigms—self-organization and global intervention strouse2022collaborating ; mosqueira2023human ; wang2025m3hf —using multi-agent influence diagrams (MAIDs), and then introduce our proposed targeted intervention paradigm.
@@ 3.2 Interpreting MARL Interaction Paradigms
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi . of MARL learning algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the self-organization paradigm are cyclic, implying computational or theoretical intricacies in finding solutions (Remark 2.5). In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependencies colas2022autotelicagentsintrinsicallymotivated . While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on targeted intervention provides a distinct advantage: it achieves solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention.
@@ 3.3.1 Definition of Pre-Strategy Intervention
To regulate the intervened agent to maximize the desired outcome $U$, we propose to add a pre-decision variable $D_{pre}$ as a new parent to the decision variable $D_{h}$ of a selected agent $h\in\mathcal{I}$. In analogy to decision variables in MAIDs, we need to assign a strategy $\sigma_{pre}$ which we refer to as pre-strategy, and this operation is named pre-strategy intervention. This definition respects the convention of stochastic intervention in causal Bayesian networks (pearl2009causality, )[Chap. 4]. A pre-strategy is determined by a pre-policy denoted by $\delta_{pre}$, which processes the agent’s information and the guidance signal, as detailed in Definition 3.2.
@@ 4 MARL for Sequential Decision Making in MAIDs
We now show how the pre-strategy intervention introduced above is implemented in MARL for sequential decision making. We first show that each Markov game can be transformed into one multi-agent influence diagram (MAID). We model MARL using team reward Markov games littman1994markov . The main idea is establishing a connection that Markov games can be transformed into MAIDs, allowing our framework to apply. The formal definition of a Markov game and the detailed process of representing it as an MAID by matching their respective variables are provided in Appendix G. Building on this foundation, the subsequent subsections will extend the MARL interaction paradigms introduced before to Markov games (Section 4.1), discuss the solvability of MARL learning paradigms based on these MARL interaction paradigms (Section 4.2), and specify how the pre-strategy intervention in Eq. (3) is integrated into generic MARL learning algorithms (Section 4.3).
@@ 4.1 Extending MARL Interaction Paradigms to Markov Games
4.2 Solvability of MARL Learning Paradigms under Various Interaction Paradigms
Independent Learning. Independent learning (IL) algorithms  [^3] Note that the IL algorithms we mention here exclude those with additional mechanisms designed to improve convergence, such as the optimistic mechanism (lauer2000algorithm, ). are the direct application of single-agent reinforcement learning algorithms, such as temporal-difference (TD) learning (sutton2018reinforcement, )[Chap. 6] and the actor-critic algorithms (konda1999actor, ). If a relevance graph is cyclic, it is not guaranteed to reach an NE by solely determining each agent’s decision variables with a generalized backward induction algorithm in analogy to independent learning, as discussed in Section 2.4. This aligns with the self-organization paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey, ). Equipping IL algorithms with the global intervention (Figures 3(b) and 4(b)) and the targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs. Such acyclic relevance graphs can guarantee the tampered IL algorithms reaching an NE with theoretical guarantees.
Centralized Training. As discussed in Remark 2.5, the only way to address cyclic relevance graphs is enabling two agents’ decision making to be matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph . More specifically, a maximal SCC consists of all the decision variables forming a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^4] A component graph is always acyclic (cormen2022introduction, ). by the generalized backward induction algorithm can reach an NE. This is in coincidence associated with the centralized training employed to reach the maximum cumulative team rewards in Markov games (oliehoek2008optimal, ; littman2001value, ). In summary, centralized training can be seen as an algorithmic approach to address the issues of pure IL algorithms existing in the self-organization paradigm, rather than switching to other interaction paradigms.
@@ 5.1 Experimental Setups
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (self-organization paradigms): for MPE, these include DQN (mnih2015human, ), VDN (sunehag2017value, ) and QMIX (rashid2020monotonic, ); for Hanabi, we use IPPO (schulman2017proximal, ), MAPPO (yu2022surprising, ) and PQN (gallici2024simplifying, ) (with PQN-IQL for independent learning and PQN-VDN for centralized training variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization examines the effectiveness of the targeted intervention paradigm without the confounding effect of the pre-policy module. This ablated variant is referred to as Intrinsic Reward that still satisfies the targeted interaction paradigm, with removing pre-decision variables $D_{pre}$ in Figures 3 and 4 and connecting the $Z^{t}$ interpreted as the intrinsic rewards to maximize directly to the decision variables $D$. To emphasize the strength of our targeted intervention paradigm, we also compare against the existing approaches implemented in the global intervention paradigm, modifying LIIR du2019liir or LAIES liu2023lazy as representative policy-based or value-based methods. This baseline is referred to as Global Intervention.
@@ 5.2 Results and Analysis
Verification of Relevance Graph Analysis. Our MAID and relevance graph analysis predicts that PSI improves the solvability of MARL learning paradigms (Section 4.2), enabling the pure independent learning more solvable. Experiments verify this across different settings. In simultaneous-move MPE (analysis in Figure 3(c) and results in Figure 5(a)), Independent Q-Learning (IQL) equipped with our PSI achieves task completion comparable to VDN, a centralized training method. This is a notable improvement for IQL over its baseline performance (without our PSI). Similarly, in sequential-move Hanabi (analysis in Figure 4(c) and results in Figure 5(c)), independent learning algorithms augmented with the PSI achieve performance comparable to, or even better than centralized training.

### v3

@@ Abstract
Steering cooperative multi-agent reinforcement learning (MARL) towards desired outcomes is challenging, particularly when the global guidance from a human on the whole multi-agent system is impractical in a large-scale MARL. On the other hand, designing external mechanisms (e.g., intrinsic rewards and human feedback) to coordinate agents mostly relies on empirical studies, lacking a easy-to-use research tool. In this work, we employ multi-agent influence diagrams (MAIDs) as a graphical framework to address the above issues. First, we introduce the concept of MARL interaction paradigms (orthogonal to MARL learning paradigms), using MAIDs to analyze and visualize both unguided self-organization and global guidance mechanisms in MARL. Then, we design a new MARL interaction paradigm, referred to as the targeted intervention paradigm that is applied to only a single targeted agent, so the problem of global guidance can be mitigated. In implementation, we introduce a causal inference technique—referred to as Pre-Strategy Intervention (PSI)—to realize the targeted intervention paradigm. Since MAIDs can be regarded as a special class of causal diagrams, a composite desired outcome that integrates the primary task goal and an additional desired outcome can be achieved by maximizing the corresponding causal effect through the PSI. Moreover, the bundled relevance graph analysis of MAIDs provides a tool to identify whether an MARL learning paradigm is workable under the design of an MARL interaction paradigm. In experiments, we demonstrate the effectiveness of our proposed targeted intervention, and verify the result of relevance graph analysis.
@@ 1 Introduction
These MARL interaction paradigms cannot be fully captured by the conventional MARL learning paradigms (i.e., independent learning (IL) (albrecht2024multi, )[Ch. 9.3] and centralized training with decentralized execution (CTDE) (albrecht2024multi, )[Ch. 9.1.3]).
@@ 2 Multi-Agent Influence Diagrams (MAIDs)
One way to address the above issue is making these two agents’ decision rules matched (koller2003multi, ), e.g. choosing both agents’ decision rules together.
@@ 3 Multi-Agent Reinforcement Learning Interaction Paradigms
We introduce the concept of MARL interaction paradigms (orthogonal to MARL learning paradigms) to describe the ways in which agents’ interactions are structured—ranging from unguided self-organization to externally guided mechanisms. Specifically, we first formalize two existing MARL interaction paradigms—self-organization and global intervention strouse2022collaborating ; mosqueira2023human ; wang2025m3hf —using multi-agent influence diagrams (MAIDs), and then introduce our proposed targeted intervention paradigm.
@@ 3.2 Interpreting MARL Interaction Paradigms
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi . of MARL learning algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the self-organization paradigm are cyclic, implying computational or theoretical intricacies in finding solutions. In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependencies colas2022autotelicagentsintrinsicallymotivated . While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on the targeted intervention paradigm provides a distinct advantage: it achieves solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention paradigm.
@@ 3.3.1 Definition of Pre-Strategy Intervention
To regulate the intervened agent to maximize the desired outcome $U$, we propose to add a pre-decision variable $D_{pre}$ as a new parent to the decision variable $D_{h}$ of a selected agent $h\in\mathcal{I}$. In analogy to decision variables in MAIDs, we need to assign a strategy $\sigma_{pre}$ which we refer to as pre-strategy, and this operation is named pre-strategy intervention. This definition respects the convention of stochastic intervention in causal Bayesian networks (pearl2009causality, )[Ch. 4]. A pre-strategy is determined by a pre-policy denoted by $\delta_{pre}$, which processes the agent’s information and the guidance signal, as detailed in Definition 3.2.
@@ 4 MARL for Sequential Decision Making in MAIDs
We now show how the pre-strategy intervention introduced above is implemented in MARL for sequential decision making. We first show that each Markov game can be transformed into one multi-agent influence diagram (MAID). We model MARL using team reward Markov games littman1994markov . The main idea is establishing a connection that Markov games can be transformed into MAIDs, allowing our framework to apply. The formal definition of a Markov game and the detailed process of representing it as an MAID by matching their respective variables are provided in Appendix G. Building on this foundation, the subsequent subsections will extend the MARL interaction paradigms introduced before to Markov games (Section 4.1), discuss the solvability of MARL learning paradigms under these MARL interaction paradigms (Section 4.2), and specify how the pre-strategy intervention in Eq. (3) is integrated into generic MARL learning algorithms (Section 4.3).
@@ 4.1 Extending MARL Interaction Paradigms to Markov Games
4.2 Solvability of MARL Learning Paradigms under MARL Interaction Paradigms
Independent Learning. Independent learning (IL) algorithms (claus1998dynamics, ) are the direct application of single-agent reinforcement learning algorithms (e.g., temporal-difference (TD) learning (sutton2018reinforcement, )[Ch. 6] and the actor-critic algorithms (konda1999actor, )) to MARL. If a relevance graph is cyclic, it is not guaranteed to reach an NE by solely determining each agent’s decision variables with a generalized backward induction algorithm in analogy to IL, as discussed in Section 2.4. This aligns with the self-organization paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey, ). Equipping IL algorithms with the global intervention (Figures 3(b) and 4(b)) and the targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs. Such acyclic relevance graphs can guarantee the tampered IL algorithms reaching an NE with theoretical guarantees.
Centralized Training and Decentralized Execution. As discussed in Remark 2.5, the only way to address cyclic relevance graphs is enabling two agents’ decision making to be matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph . More specifically, a maximal SCC consists of all the decision variables forming a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^3] A component graph is always acyclic (cormen2022introduction, ). by the generalized backward induction algorithm can reach an NE. This is in coincidence associated with the centralized training and decentralized execution (CTDE) employed to reach the maximum cumulative team rewards in Markov games (oliehoek2008optimal, ; littman2001value, ). In summary, CTDE can be seen as an MARL learning paradigm to address the non-stationarity dilemma of vanilla IL algorithms existing in the self-organization paradigm, rather than switching to other MARL interaction paradigms.
@@ 5.1 Experimental Setups
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (self-organization paradigms): for MPE, these include DQN (mnih2015human, ), VDN (sunehag2017value, ) and QMIX (rashid2020monotonic, ); for Hanabi, we use IPPO (schulman2017proximal, ), MAPPO (yu2022surprising, ) and PQN (gallici2024simplifying, ) (with PQN-IQL for IL and PQN-VDN for CTDE variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization examines the effectiveness of the targeted intervention paradigm without the confounding effect of the pre-policy module. This ablated variant is referred to as Intrinsic Reward that still satisfies the targeted interaction paradigm, with removing pre-decision variables $D_{pre}$ in Figures 3 and 4 and connecting the $Z^{t}$ interpreted as the intrinsic rewards to maximize directly to the decision variables $D$. To emphasize the strength of our targeted intervention paradigm, we also compare against the existing approaches implemented in the global intervention paradigm, modifying LIIR du2019liir or LAIES liu2023lazy as representative policy-based or value-based methods. This baseline is referred to as Global Intervention.
@@ 5.2 Results and Analysis
Verification of Relevance Graph Analysis. Our MAID and relevance graph analysis predicts that PSI improves the solvability of MARL learning paradigms (Section 4.2), enabling the vanilla IL learning paradigm more solvable. Experiments verify this across different settings. In simultaneous-move MPE (analysis in Figure 3(c) and results in Figure 5(a)), Independent Q-Learning (IQL) equipped with our PSI achieves task completion comparable to VDN, a CTDE algorithm. This is a notable improvement for IQL over its baseline performance (without our PSI). Similarly, in sequential-move Hanabi (analysis in Figure 4(c) and results in Figure 5(c)), IL algorithms augmented with the PSI achieve performance comparable to, or even better than CTDE algorithms.